OpenStudy (kalebj):

What is (1.9 x 10 to the 3rd power) x (2.3 x 10 to the 4th power) ? A.4.2 x 10 to the 7th power B.4.37 x 10 to the 7th power C.4.2 x 10 to the 12th power D.4.37 x 10 to the 12th power

OpenStudy (kalebj):

@KJSaif

OpenStudy (kalebj):

@mathmate

OpenStudy (kalebj):

@Ms-Brains

OpenStudy (kalebj):

I have no idea what to do at all

OpenStudy (mathmate):

Use rules of exponents: \((a\times 10^x)(b\times 10^y)=ab\times 10^{x+y}\)